The length of a side of a square is 5x+2 units.does it have a perimeter of 48 if it os what is equation for it? What is the x? What is the area of the square?

Answer by algebrahouse.com(1659)

To find the perimeter of a square, take 4 times the length of one side:

4(5x + 2) = 48 {four times one side equals the possible perimeter}
20x + 8 = 48 {used distributive property}
20x = 40 {subtracted 8 from each side}
x = 2 {divided each side by 0}

To find the area of a square, just square one side:
5x + 2 is one side
= 5(2) + 2 {substituted 2 for x}
= 10 + 2 {multiplied}
= 12 {added}

therefore, one side is 12
the area is 12², which is 144 square units

For more help from me, visit: www.algebrahouse.com